I REALLY NEED HELP WITH THESE PROBLEMS, ITS DUE TOMORROW AND I CANT SEEM TO GET THE ANSWERS TO THESE PROBELMS..IT IS ABOUT VERBAL PROBLEMS..THANKS!
THREE LESS THAN FIVE TIMES A CERTAIN NUMBER IS EQUAL TO 67 MORE THAN ONE-THIRD OF THE NUMBER. FIND THE NUMBER.
THANKS AGAIN!

Answer by longjonsilver(2297)

let the number be x

so, one side is 5x-3
and other side is x/3 + 67

So, 5x-3 = x/3 + 67
5x = x/3 + 70
15x = x + 210
14x = 210
x = 210/14
x = 15

jon.